# Exploiting the Final Component of Generator Architectures for AI-Generated Image Detection

### v1

@@ 1 Introduction
When examining objects like a careful detective, the last touch often reveals the most, having overwritten the layers of history beneath it. Similarly, given an image and the task of determining whether it is real or AI-generated, it is crucial to examine the distinctive traces left by the generator. Contemporary image generators are no longer monolithic; rather, they comprise multiple modular components operating in stages. This raises a natural and important question: does the last component, which serves as the final step in producing the image, leave identifiable evidence that can be exploited for detection? Our intuition is that, regardless of the generation paradigm—diffusion-based [13], autoregressive [27], or otherwise—there inevitably exists a concluding step that transforms intermediate representations into pixels, serving as the generator’s “final touch.”
@@
We approach this question from the perspective of architecture. To clarify, the generation architecture refers to network components and how they are interconnected. A network component is a functional block composed of multiple neural layers, such as a VAE decoder, DiT transformer, or tokenizer. In contrast, most existing deepfake detection approaches focus on the generation model, defined as a specific set of trained weights (e.g., a pretrained network and its fine-tuned variant are considered different models). Current approaches typically generate training samples from specific generation models and design sophisticated features—such as CLIP embeddings [44], frequency-domain cues [28], multi-scale representations [19], reverse-engineered traces [16], or gradient-based signals [10]—to distinguish synthetic images from real ones. However, as new generators continue to emerge, the generalizability of existing detection approaches becomes a critical concern. When the distribution of testing data shifts significantly, these approaches often require regenerating of synthetic images from new models in order to retrain the detectors. The challenge is further amplified when users fine-tune models on custom data and keep them private.
@@
Since real images are readily available from public datasets for image-related tasks (e.g., ImageNet, MS-COCO), we propose to “contaminate” real images with generation traces using only the generator’s final component, and subsequently train a detector to differentiate them from the original real images. Operating samples through a single component is substantially faster than executing the entire generation pipeline, particularly in diffusion models that require iterative multi-step inference. Moreover, this strategy does not necessitate the generator to be open-sourced; gray-box access to its final component suffices for implementation.
@@
The potential generalizability of the proposed approach is motivated by three key observations: (a) many generators share similar or even identical final components; (b) relying on only a small fraction of the total parameters reduces overfitting to specific generators; (c) the contaminated images preserve identical semantic content to their real counterparts, ensuring that the detector focuses on generation artifacts rather than semantic information specific to the training set.
@@
To comprehensively validate our hypothesis regarding the generalizability of the generator’s final component, we propose a novel taxonomy that categorizes image generators according to their final architectural component and analyze 21 widely used generators under this framework (see Fig.1). Traditionally, image generators are grouped by generation paradigm into three major categories [40]: Generative Adversarial Networks, Diffusion models, and more recently, Autoregressive models. However, many recent generators frequently reuse common components based on their function, such as the VAE decoder, which decodes images from latent space.
@@
From our taxonomy, we identify three representative final components—VAE decoder, super-resolution diffuser, and VQ de-tokenizer—and employ them to imprint traces onto real images as part of our approach. We observe that these components correspond to distinct trace spaces. We conduct a series of experiments to investigate: (a) whether the final component alone is sufficient to detect images generated by its corresponding full architecture; (b) whether a detector trained on one component generalizes to images generated by architectures from other categories; and (c) whether the detector remains effective on fine-tuned variants of the same component. Finally, we select 100 samples from each of the three components using K-medoids clustering, combining the 300 samples to fine-tune a detector built upon a pretrained DINOv3 [32] backbone with an appended fully-connected layer. The resulting detector achieves strong zero-shot performance across multiple challenging benchmarks.
@@ 2.1 Architecture attribution
For the deepfake detection task that this study focuses on, exploration from the perspective of generative architectures remains limited. In the related task of synthesized image attribution, which aims to identify the generator responsible for a given image, few studies have investigated attribution at the architectural level. Yang et al. [43] claimed to be the first to explore architecture attribution. They assessed GAN architectures and found that architectures leave globally consistent fingerprints, whereas model weights produce regional traces. [41] and [42] reveal that similar diffusion architectures produce similar patterns in frequency-domain features and achieve similar detection accuracies. [1] proposed a Siamese network to verify whether two images originate from the same generation architecture, while [31] applied one-class classification for open-set architecture scenarios.
@@ 2.3 Zero-shot deepfake detection
3 Method
We aim to investigate whether the final component leaves identifiable traces in generated images when all preceding components are omitted, and how well such traces generalize across different generation architectures. The core idea is to use real images to elicit generation signatures and train a binary classifier on them. Let the training set of real images be denoted as $\mathcal{D}=\{\mathbf{x}\}$, where each image $\mathbf{x}\in\mathbb{R}^{h\times w\times 3}$ is labeled as 1 (indicating a real image). Let $\varphi^{*}(\cdot):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$ represent the final component of a well-trained generative model, where $\mathbb{S}^{d}$ is an internal space. To simplify notation, we do not distinguish the model’s parameter set from its corresponding mapping function, e.g., $\varphi^{*}$ denotes both the mapping function implemented by a network module and its learnable parameters. The objective function of the proposed AI-generated image detection is the standard cross-entropy on real images and reconstructed counterparts, as in Eq. 1:
$\mathcal{L}(\theta;\mathcal{E}):=-E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta(\mathbf{x})+log(1-\theta(\varphi^{*}(\mathcal{E}(\mathbf{x})))]$ (1)
where $\mathcal{E}(\cdot):\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{S}^{d}$ is a function that maps RGB image into internal space. In generative models, $\varphi^{*}$ is optimized from the objective of ELBO:
$(\varphi^{*},\mathcal{E}^{*})=$ $\arg\min_{\varphi,\mathcal{E}}\Bigl(-E_{\mathbf{z}\sim q_{\mathcal{E}}(\mathbf{z}|\mathbf{x})}[\log(p_{\varphi}(\mathbf{x}|\mathbf{z}))]+\bigr$.
$\bigl.D_{KL}(q_{\mathcal{E}}(\mathbf{z}|\mathbf{x})||p(\mathbf{z}))\Bigr)$ (2)
Given a test image $\mathbf{x}_{t}$, the classifier trained with Eq. 1 outputs $\theta(\mathbf{x}_{t})$, indicating whetherr $\mathbf{x}_{t}$ is real ($1$) or AI-generated (0).
Property 1 If $\mathcal{E}^{*}$ is used as the mapping function $\mathcal{E}$ in Eq.1, then the expected detection error for images generated via $\varphi^{*}$ is not greater than that obtained using any other mapping function $\mathcal{E}$, i.e., $E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta_{\mathcal{E}^{*}}(\mathbf{x})]-E_{\mathbf{z}\sim q_{\mathcal{E}^{*}}(\mathbf{z}|\mathbf{x})}[log(1-\theta_{\mathcal{E}^{*}}(\varphi^{*}(\mathbf{z}))]\leq-E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta_{\mathcal{E}}(\mathbf{x})]-E_{\mathbf{z}\sim q_{\mathcal{E}^{*}}(\mathbf{z}|\mathbf{x})}[log(1-\theta_{\mathcal{E}}(\varphi^{*}(\mathbf{z}))]$, where $\theta_{\mathcal{E}}=\arg\min\mathcal{L}(\cdot;\mathcal{E})$ and $\theta_{\mathcal{E}^{*}}=\arg\min\mathcal{L}(\cdot;\mathcal{E}^{*})$.
This is an idealized property concerning global minimizers, yet it suggests that using the optimal mapping function $\mathcal{E}^{*}$, which was jointly trained with $\varphi^{*}$ in the original generator, results in a lower expected detection error.
[…]
$\hat{\nabla}_{\theta}^{\text{paired}}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\Big[\nabla_{\theta}\log\theta(\mathbf{x})+\nabla_{\theta}\log\big(1-\theta(\varphi(\mathcal{E}(\mathbf{x})))\big)\Big]$
$\hat{\nabla}_{\theta}^{\text{indep}}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\nabla_{\theta}\log\theta(\mathbf{x})-\frac{1}{B}\sum_{\mathbf{x}^{\prime}\in\mathcal{B}^{\prime}_{t}}\nabla_{\theta}\log\big(1-\theta(\mathbf{x}^{\prime})\big),$
where $\mathcal{B}_{t}\subseteq\mathcal{D}$ and $\mathcal{B}^{\prime}_{t}\subseteq\mathcal{D}^{\prime}$ are mini-batches of size $B$.
A theoretical and empirical discussion is provided in Appendix. This property suggests that independently shuffling real and constructed samples reduces the variance of the stochastic gradient, which in turn leads to more stable and faster convergence in SGD [29, 3]. Therefore, it is advantageous to construct the negative sample set $\mathcal{D}^{\prime}$ offline and sample from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ independently during SGD training.
@@ 3 Method
Fig.2 presents an overview of our method. We next detail the two key components of the objective function Eq.1: the fake-sample construction $\varphi\circ\mathcal{E}(\cdot)$ and the detector $\theta$. Before delving into these, we examine the final components employed by popular text-to-image generators and propose a taxonomy to systematically analyze their architectural design choices.
3.1 A Taxonomy of Generation Architectures based on Final Component
The design principles of our taxonomy, grounded in the last component $\varphi(\mathbf{z}):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$, are as follows: (1) Position—Last: The output of $\varphi(\mathbf{z})$ must be an image in RGB space $\mathbb{R}^{h\times w\times 3}$. We define the taxonomy categories based on the input $\mathbf{z}$ and its input space $\mathbb{S}^{d}$. (2) Granularity—Middle-level: Deep network units can be considered at four levels of granularity. The finest level refers to individual layers, such as activation or pooling layers. The second level refers to blocks composed of multiple layers that serve highly detailed functions, such as upsampler in U-Net or transformer heads. The coarsest level is the entire model. “Component” lies in the third level, representing a functional module that operates independently, such as VAE decoders or VQ de-tokenizer. This choice is motivated by two factors: many generators release pre-trained weights for such components separately, making them more accessible without the full pipeline; and fine-grained units, like individual layers or small blocks, leave traces that are too subtle or indistinct for reliable detection. (3) Perspective—Implementation. The taxonomy is based on functional implementation rather than the high-level role of the component. For example, MAR [20] performs autoregressive prediction over masked tokens but ends with a diffusion denoiser, so it is categorized as “denoiser” rather than “de-tokenizer” like other autoregressive generators. DeepFloyd IF [9] performs super-resolution in latent space and ends with a VAE decoder, placing it in the decoder category rather than denoiser like other super-resolution generators.
@@ 3.1 A Taxonomy of Generation Architectures based on Final Component
As illustrated in Fig.1, we define four primary categories based on the input space $\mathbb{S}^{d}$, and further subdivide them according to what $\mathbf{z}$ represents:
@@
  • VAE.decoder. In this category, the final component operates in a continuous latent space, i.e., $\mathbb{S}^{d}:=\mathbb{R}^{d}$, where $d\ll h\times w\times 3$. It decodes a $d$-dimensional feature map or vector into the final image. As many modern generators target high-resolution outputs (e.g., $2048\times 2048$), they often perform generation in a low-dimensional latent space for computational efficiency, relying on encoder–decoder structures for dimensionality reduction and expansion. There are three typical functional contexts in which $\mathbf{z}$ resides: (1) Latent diffusion, where diffusion models operate directly in the latent space, as in the Stable Diffusion family [34] and FLUX [5]. As illustrated in Fig.1(a), in this case, $\mathbf{z}$ represents the denoised latent sample predicted by the diffusion model. (2) Enhancement, where the latent representation from the diffusion model is further refined—e.g., via GAN-based distillation—to produce a more precise $\mathbf{z}$ before decoding, as in HiDream [4]. (3) Latent super-resolution, where a low-resolution image is first generated and then progressively upscaled through latent-space super-resolution stages. As shown in Fig.1(b), the low-resolution image is noised, resized to the target size, and encoded into a latent space for diffusion processing. The resulting upscaled latent, $\mathbf{z}$, is then decoded to produce the final high-resolution image. Examples include CogView 4 [45] and DeepFloyd IF [9], which employ the Stable Diffusion 4× Upscaler [30] as the final super-resolution stage.
    Algorithm 1 Negative sample construction from VAE_1.1
@@
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: vae=AutoencoderKL.from_pretrained(“sd-2-1”, subfolder=“vae”)
@@
    Algorithm 2 Negative sample construction from VQ_2.2
@@
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: vq=AutoModelForCausalLM.from_pretrained(“janus”).vq_model
@@
    3:  $\mathbf{z}$ = vq.quantize(vq.encode($\mathbf{x}$))
    4:  $\hat{\mathbf{x}}$ = vq.decode($\mathbf{z}$)
    5:end for
    6:return $\{\hat{\mathbf{x}}\}$
    Algorithm 3 Negative sample construction from Super-resolution_3.3
    Input: A real image set $\mathcal{D}=\{\mathbf{x}\}$
    Output: The constructed image set $\{\hat{\mathbf{x}}\}$
    1: denoiser=PixelFlow(“model.pt”)
@@
  • VQ.de-tokenizer. The final component operates in discrete token space, i.e., $\mathbb{S}^{d}:=\mathbb{Q}^{p}$, where $\mathbb{Q}$ is a learned codebook and $p$ denotes the number of spatial tokens (i.e., divided image patches) in the quantized representation. There are three common contexts for $\mathbf{z}$: (1) Next-token auto-regression, where tokens are predicted sequentially and the VQ decoder transforms them back into image pixels, as in Emu 3 [38], JanusPro [8], and LlamaGen [35]. (2) Next-scale autoregression, where the generator starts from coarse-scale tokens (e.g., a $1\times 1$ token map) and progressively predicts higher-resolution token maps conditioned on all previous ones, as in VAR [36] and Infinity [12]. (3) Latent diffusion hybrid, which combines global latent diffusion for coarse image structure with autoregressive token prediction for sequentially refining discrete latent components, e.g., Kandinsky [2].
  • Diffusion.denoiser. In this category, the final step is denoising, applied over three typical input spaces. (1) Continuous token: $\mathbb{S}^{d}:=\mathbb{R}^{p}$. The image is divided into $p$ patches, with tokens represented as continuous embeddings rather than discrete codes. MAR [20] exemplifies this approach, as illustrated in Fig.1(d). (2) Low-resolution image: $\mathbb{S}^{d}:=\mathbb{R}^{h_{l}\times w_{l}\times 3}$, where $h_{l}<h,w_{l}<w$. In this setting, the generator builds images progressively from low to high resolution through multiple upscaling stages, e.g., $64\times 64\rightarrow 256\times 256\rightarrow 1024\times 1024$. The final step adds noise to the previous resolution’s image, resizes it to the target size, and applies diffusion denoising. This process is similar to VAE.decoder–latent super-resolution but operates directly in pixel space (see Fig.1(e)). Two training strategies are common: cascade super-resolution, where each adjacent low-to-high resolution stage is trained separately (e.g., DALL·E 3 [26] and GLIDE [24]), and end-to-end super-resolution, where all stages are trained jointly (e.g., PixelFlow [7]). (3) Same-size pixels: $\mathbb{S}^{d}:=\mathbb{R}^{h\times w\times 3}$. The input $\mathbf{z}$ has the same spatial resolution as the target image. It can either represent a low-quality image to be refined by the diffusion denoiser, as in ReflectionFlow [47], or an initial Gaussian noise sample used in conventional diffusion models such as DDPM [13] and DDIM [33] (Fig.1(f)).
@@
  • One-stop.generator. In this setting, $\mathbb{S}^{d}:=\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. The generator directly maps a sampled noise vector to the final image, as shown in Fig.1(g). This category includes GAN-based approaches such as GigaGAN [14], which produce high-resolution images through large-scale adversarial training, and pixel-wise auto-regressive models like JetFormer [37], which generate images sequentially in an end-to-end fashion.
The categorization of these representative generators in our proposed taxonomy is based on their official source code or released documentation, with detailed references provided in the Appendix.

### v2

@@ 1 Introduction
When examining objects like a careful detective, the last touch often reveals the most, having overwritten the layers of history beneath it. Likewise, when determining whether an image is real or AI-generated, the key is to examine the distinctive traces left by the generator, partcularly those from its final stage. Contemporary image generators are no longer monolithic; rather, they comprise multiple modular components operating in stages. This raises a natural and important question: does the last component, which serves as the final step in producing the image, leave identifiable evidence that can be exploited for detection? Our intuition is that, regardless of the generation paradigm—generative adversarial[16], diffusion-based[19], autoregressive[34], or otherwise—there inevitably exists a concluding step that transforms intermediate representations into pixels, serving as the generator’s “final touch”.
@@
We approach this question from the perspective of generation architecture. To clarify, the architecture refers to the structural design of a network, including its functional components and their interconnections. A network component is a functional module composed of multiple neural layers, such as a VAE decoder, DiT transformer, or tokenizer. In contrast, the generation model denotes a specific instantiation of trained weights (e.g., a pretrained network and its fine-tuned variant are considered different models). Most existing deepfake detection methods operate at the model level. They typically generate training samples from particular generation models and design sophisticated features—such as CLIP embeddings[51], frequency-domain cues [35], multi-scale representations [25], reverse-engineered traces [22], or gradient-based signals [14]—to distinguish synthetic images from real ones. However, as new generators continue to emerge, the generalizability of these methods becomes a critical concern. When the distribution of testing data shifts significantly, these approaches often require regenerating of synthetic images from new generation models to retrain the detectors. This challenge is further amplified when users fine-tune generation models on custom data and keep the resulting models private.
@@
Since real images are readily available from public datasets for image-related tasks (e.g., ImageNet [13], MS-COCO [28]), we propose to “contaminate” real images with generation traces using only the generator’s final component, and subsequently train a detector to differentiate them from the original real images. Operating samples through a single component is substantially faster than executing the entire generation pipeline, particularly in diffusion models that require iterative multi-step inference. Moreover, this strategy does not necessitate the generator to be open-sourced; gray-box access to its final component suffices for implementation.
@@
The potential generalizability of the proposed approach is motivated by three key observations: (a) many generators share similar or even identical final components; (b) relying on only a small fraction of the total parameters reduces overfitting to specific generators; (c) the contaminated images preserve identical semantic content to their real counterparts, ensuring that the detector focuses on generation artifacts rather than semantic information.
@@
To comprehensively validate our hypothesis regarding the generalizability of the generator’s final component, we propose a novel taxonomy that categorizes image generators according to their final architectural component and analyze 21 widely used generators under this framework. Traditionally, image generators are grouped by generation paradigm into three major categories [47]: Generative Adversarial Networks, Diffusion Models, and more recently, Autoregressive Models. However, many recent generators frequently reuse common components based on their function, such as the VAE decoder, which decodes images from latent space.
@@
From our taxonomy, we identify three representative final components—VAE decoder, super-resolution diffuser, and VQ de-tokenizer—and employ them to imprint traces onto real images as part of our approach. We conduct a series of experiments on them to investigate: whether the final component alone is sufficient to detect images generated by its corresponding full architecture; whether a detector trained on one component generalizes to images generated by architectures from other categories; and whether the detector remains effective on fine-tuned variants of the same component. Finally, we select 100 samples from each of the three components using K-medoids clustering, combining the 300 samples to fine-tune a detector built upon a pretrained DINOv3 [39] backbone with an appended fully-connected layer. The resulting detector achieves strong zero-shot performance across multiple challenging benchmarks.
@@ 2.1 Architecture attribution
For the deepfake detection task that this study focuses on, exploration from the perspective of generative architectures remains limited. While for the related task of synthesized image attribution, which aims to identify the generator responsible for a given image, few studies have investigated attribution at the architectural level. Yang et al. [50] claimed to be the first to explore architecture attribution. They assessed GAN architectures and found that architectures leave globally consistent fingerprints, whereas models produce regional traces. Wißmann et al. [48] and Xu et al. [49] reveal that similar diffusion architectures produce similar patterns in frequency-domain features. Furthermore, Abady et al. [1] proposed a Siamese network to verify whether two images originate from the same generation architecture, while Shahid et al. [38] applied one-class classification for open-set architecture scenarios.
@@ 2.3 Zero-shot deepfake detection
3 A Taxonomy of Generation Architectures based on Final Component
To facilitate the exploration of common final components in modern generation architectures and the generalizable traces they leave behind, we propose a novel taxonomy that categorizes image generators according to their final architectural component. Let $\varphi(\mathbf{z}):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$ denote this final component, where $\mathbb{S}^{d}$ represents the internal latent space. The design principles of our taxonomy, grounded in $\varphi(\mathbf{z})$, are as follows:
  • Position—Last: The output of $\varphi(\mathbf{z})$ must be an image in RGB space $\mathbb{R}^{h\times w\times 3}$. We define the taxonomy categories based on the input $\mathbf{z}$ and its input space $\mathbb{S}^{d}$.
  • Granularity—Middle-level: Deep network units can be considered at four levels of granularity. The finest level refers to individual layers, such as activation or pooling layers. The second level refers to blocks composed of multiple layers that serve highly detailed functions, such as upsampler in U-Net or transformer heads. The coarsest level is the entire model. “Component” lies in the third level, representing a functional module that operates independently, such as VAE decoders or VQ de-tokenizer. This choice is motivated by two factors: many generators release pre-trained weights for such components separately, making them more accessible without the full pipeline; and fine-grained units, like individual layers or small blocks, leave traces that are too subtle or indistinct for reliable detection.
  • Perspective—Implementation. The taxonomy is based on functional implementation rather than the high-level role of the component. For example, as MAR [26] performs autoregressive prediction over masked tokens but ends with a diffusion denoiser, it is categorized as “denoiser” rather than “de-tokenizer” like other autoregressive generators. DeepFloyd IF [12] performs super-resolution in latent space and ends with a VAE decoder, placing it in the decoder category rather than denoiser like other super-resolution generators.
@@ 3 A Taxonomy of Generation Architectures based on Final Component
As illustrated in Fig.1, we define four primary categories based on the input space $\mathbb{S}^{d}$, and further subdivide them according to what $\mathbf{z}$ represents, as discussed in Sections 3.1-3.4. The categorization of all representative generators in our proposed taxonomy is based on their official source code or released documentation, with detailed references provided in the Appendix.
3.1 VAE.decoder
In the category of VAE.decoder, the final component $\phi(\mathbf{z})$ operates in a continuous latent space, i.e., $\mathbb{S}^{d}:=\mathbb{R}^{d}$, where $d\ll h\times w\times 3$. It decodes a $d$-dimensional feature map or vector into the final image. As many modern generators target high-resolution outputs (e.g., $2048\times 2048$), they often perform generation in a low-dimensional latent space for computational efficiency, relying on encoder–decoder structures for dimensionality reduction and expansion.
@@
There are three typical functional contexts in which $\mathbf{z}$ resides: (1) Latent diffusion, where diffusion models operate directly in the latent space, as in the Stable Diffusion family [41] and FLUX [7]. As illustrated in Fig.1(a), in this case, $\mathbf{z}$ represents the denoised latent sample predicted by the diffusion model. (2) Enhancement, where the latent representation from the diffusion model is further refined—e.g., via GAN-based distillation—to produce a more precise $\mathbf{z}$ before decoding, as in HiDream [5]. (3) Latent super-resolution, where a low-resolution image is first generated and then progressively upscaled through latent-space super-resolution stages. As shown in Fig.1(b), the low-resolution image is noised, resized to the target size, and encoded into a latent space for diffusion processing. The resulting upscaled latent, $\mathbf{z}$, is then decoded to produce the final high-resolution image. Examples include CogView 4 [52] and DeepFloyd IF [12], which employ the Stable Diffusion 4×Upscaler [37] as the final super-resolution stage.
3.2 VQ.de-tokenizer
The final component of VQ.detokenier generators operates in discrete token space, i.e., $\mathbb{S}^{d}:=\mathbb{Q}^{p}$, where $\mathbb{Q}$ is a learned codebook and $p$ denotes the number of spatial tokens (i.e., divided image patches) in the quantized representation.
There are three common contexts for $\mathbf{z}$: (1) Next-token auto-regression, where tokens are predicted sequentially and the VQ decoder transforms them back into image pixels, as in Emu 3 [45], JanusPro [10], and LlamaGen [42]. (2) Next-scale autoregression, where the generator starts from coarse-scale tokens (e.g., a $1\times 1$ token map) and progressively predicts higher-resolution token maps conditioned on all previous ones, as in VAR [43] and Infinity [18]. (3) Latent diffusion hybrid, which combines global latent diffusion for coarse image structure with autoregressive token prediction for sequentially refining discrete latent components, e.g., Kandinsky [2].
3.3 Diffusion.denoiser
In the category of Deiffusion.denoiser, the final step is denoising, applied over three typical input spaces. (1) Continuous token: $\mathbb{S}^{d}:=\mathbb{R}^{p}$. The image is divided into $p$ patches, with tokens represented as continuous embeddings rather than discrete codes. MAR [26] exemplifies this approach, as illustrated in Fig.1(d). (2) Low-resolution image: $\mathbb{S}^{d}:=\mathbb{R}^{h_{l}\times w_{l}\times 3}$, where $h_{l}<h,w_{l}<w$. In this setting, the generator builds images progressively from low to high resolution through multiple upscaling stages, e.g., $64\times 64\rightarrow 256\times 256\rightarrow 1024\times 1024$. The final step adds noise to the previous resolution’s image, resizes it to the target size, and applies diffusion denoising. This process is similar to VAE.decoder–latent super-resolution but operates directly in pixel space (see Fig.1(e)). Two training strategies are common: cascade super-resolution, where each adjacent low-to-high resolution stage is trained separately (e.g., DALL·E 3 [33] and GLIDE [31]), and end-to-end super-resolution, where all stages are trained jointly (e.g., PixelFlow [9]). (3) Same-size pixels: $\mathbb{S}^{d}:=\mathbb{R}^{h\times w\times 3}$. The input $\mathbf{z}$ has the same spatial resolution as the target image. It can either represent a low-quality image to be refined by the diffusion denoiser, as in ReflectionFlow [54], or an initial Gaussian noise sample used in conventional diffusion models such as DDPM [19] and DDIM [40] (Fig.1(f)).
Algorithm 1 Sample construction from VAE
@@
Output: The constructed
image set $\{\hat{\mathbf{x}}\}$
1:  Load vaevae model
@@
Algorithm 2 Sample construction from VQ
@@
Output: The constructed
image set $\{\hat{\mathbf{x}}\}$
1:  Load vqvq model
@@
3:   $h$ =vq.encode($\mathbf{x}$)
4:   $\mathbf{z}$ = vq.quantize($h$)
5:   $\hat{\mathbf{x}}$ = vq.decode($\mathbf{z}$)
6:  end for
7:  return $\{\hat{\mathbf{x}}\}$
Algorithm 3 Sample construction from Diffusion
Input: Real images $\mathcal{D}=\{\mathbf{x}\}$
Output: Constructed set $\{\hat{\mathbf{x}}\}$
1:  Load denoiserdenoiser model
@@
Algorithm 4 Sparse sample selection
Input: Constructed image sets: $\mathcal{D}_{1}\cup\mathcal{D}_{2}\cup\mathcal{D}_{3}=\{\hat{\mathbf{x}}\}_{\text{VAE}}\cup$ $\{\hat{\mathbf{x}}\}_{\text{VQ}}\cup\{\hat{\mathbf{x}}\}_{\text{Diffusion}}$, the number of samples to be selected: $k$
Output: The sparse constructed image set $\{\hat{\mathbf{x}}\}$
1:  Load pretrained dinov3 model
2:  for $i=1$ to 33 do
4:   dist = Euclidean_distances(features)
5:   $\mathcal{M}_{i}$= Kmedoids(dist, $k$)      //a set of $k$ instances of $\hat{\mathbf{x}}$
6:  end for
7:  return $\mathcal{M}_{1}\cup\mathcal{M}_{2}\cup\mathcal{M}_{3}$
3.4 SingleStage.generator
@@
In the setting of SingeStage.generator, $\mathbb{S}^{d}:=\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$. The generator directly maps a sampled noise vector to the final image, as shown in Fig.1(g). This category includes GAN-based approaches such as GigaGAN [20], which produce high-resolution images through large-scale adversarial training, and pixel-wise auto-regressive models like JetFormer [44], which generate images sequentially in an end-to-end fashion.
4 Method
We aim to investigate whether the final component leaves identifiable traces in generated images when all preceding components are omitted, and how well such traces generalize across different generation architectures. The core idea is to use real images to elicit generation signatures and train a binary classifier on them. Let the training set of real images be denoted as $\mathcal{D}=\{\mathbf{x}\}$, where each image $\mathbf{x}\in\mathbb{R}^{h\times w\times 3}$ is labeled as 1 (indicating a real image). Let $\varphi^{*}(\cdot):\mathbb{S}^{d}\rightarrow\mathbb{R}^{h\times w\times 3}$ represent the optimized final component of a well-trained generative model, where $\mathbb{S}^{d}$ is an internal space. To simplify notation, we do not distinguish the model’s parameter set from its corresponding mapping function, e.g., $\varphi^{*}$ denotes both the mapping function implemented by a network module and its learnable parameters.
The objective function of the proposed AI-generated image detection is the standard cross-entropy on real images and reconstructed counterparts:
$\mathcal{L}(\theta;\mathcal{E}):=-E_{\mathbf{x}\sim p_{\mathcal{D}}}[log(\theta(\mathbf{x}))+log(1-\theta(\varphi^{*}(\mathcal{E}(\mathbf{x})))]$ (1)
where $\mathcal{E}(\cdot):\mathbb{R}^{h\times w\times 3}\rightarrow\mathbb{S}^{d}$ maps an RGB image into the latent space. Given a test image $\mathbf{x}_{t}$, the classifier trained with Eq. 1 outputs $\theta(\mathbf{x}_{t})$, indicating the probability that $\mathbf{x}_{t}$ is real ($1$) or AI-generated (0). In this formulation, The encoder $\mathcal{E}$ determines the representation from which the corresponding fake image $\mathbf{x}$ is decoded. It is predefined and not co-trained with $\theta$. The following property formulates the intuition behind the choice of $\mathcal{E}$.
Property 1 (Choice of encoder) Let $\mathcal{E}^{*}$ be the encoder jointly optimized with $\varphi^{*}$ during generator training, i.e., $(\mathcal{E}^{*},\varphi^{*})=\arg\min\mathcal{L}_{g}(\mathcal{E},\varphi)$, where $\mathcal{L}_{g}$ is the generation loss. For any encoder $\mathcal{E}$, under global optimality, $\mathcal{L}(\theta_{\mathcal{E}^{*}},\mathcal{E}^{*})\leq\mathcal{L}(\theta_{\mathcal{E}},\mathcal{E})$, where $\theta_{\mathcal{E}}=\arg\min_{\theta}\mathcal{L}(\theta;\mathcal{E})$.
This property is stated under the idealized assumption of global minimizers. Nevertheless, it suggests that the encoder $\mathcal{E}^{*}$, being jointly optimized and distributionally aligned with the decoder $\varphi^{*}$, achieves the minimal attainable expected detection loss among all possible mapping functions $\mathcal{E}$.
Figure 3: Feature visualization of the three final components.
Figure 4: Sparse sample selection.
Property 2 (Non-pairwise training) Let $\mathcal{D}=\{\mathbf{x}\}$ be real samples and $\mathcal{D}^{\prime}=\{\mathbf{x}^{\prime}$ $\varphi(\mathcal{E}(\mathbf{x}))\mid\mathbf{x}\in\mathcal{D}\}$ be constructed samples. Sampling mini-batches independently from $\mathcal{D}$ and $\mathcal{D}^{\prime}$, rather than using paired $(\mathbf{x}_{i},\mathbf{x}^{\prime}_{i})$, yields a gradient estimator with lower variance by SGD training, i.e., $\text{Var}(\hat{\nabla}_{\theta}^{(I)})\leq\text{Var}(\hat{\nabla}_{\theta}^{(P)})$, where
$\hat{\nabla}_{\theta}^{(I)}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\nabla_{\theta}\log\theta(\mathbf{x})-\frac{1}{B}\sum_{\mathbf{x}^{\prime}\in\mathcal{B}^{\prime}_{t}}\nabla_{\theta}\log\big(1-\theta(\mathbf{x}^{\prime})\big),$
$\hat{\nabla}_{\theta}^{(P)}=-\frac{1}{B}\sum_{\mathbf{x}\in\mathcal{B}_{t}}\Big[\nabla_{\theta}\log\theta(\mathbf{x})+\nabla_{\theta}\log\big(1-\theta(\varphi(\mathcal{E}(\mathbf{x})))\big)\Big],$
and $\mathcal{B}_{t}\subseteq\mathcal{D}$, $\mathcal{B}^{\prime}_{t}\subseteq\mathcal{D}^{\prime}$ are mini-batches of size $B$.
A theoretical and empirical justification is provided in Appendix. This property suggests that independently shuffling real and constructed samples reduces the variance of the stochastic gradient, which in turn leads to more stable and faster convergence in SGD [36, 4]. Therefore, it is advantageous to construct the negative sample set $\mathcal{D}^{\prime}$ offline and sample from $\mathcal{D}$ and $\mathcal{D}^{\prime}$ independently during SGD training.
Fig. 2 presents an overview of our method. The input is a real image, and the corresponding fake samples are constructed from it using the encoder and decoder. This operation realizes $\varphi\circ\mathcal{E}(\cdot)$ in Eq. 1, leaving detectable traces while preserving the exact semantic content of the real image. The detector $\theta$ is implemented using a pretrained feature extractor (DINOv3) with an additional classification layer. In the following, we detail these two key components of the objective function in Eq. 1: the fake sample construction $\varphi\circ\mathcal{E}(\cdot)$ and the detector $\theta$.